OpenStudy (anonymous):

indefinite integral of t^2 (t+2)^(1/2)

OpenStudy (idku):

you can do integration by parts, or add a magic zero a munch of times.

OpenStudy (idku):

By magic zero I mean, \[\int\limits_{ }^{ }t^2\sqrt{t+2}+(4t+4)\sqrt{t+2}-(4t+4)\sqrt{t+2}dt\]\[\int\limits_{ }^{ }(t^2+4t+4)\sqrt{t+2}-(4t+4)\sqrt{t+2}dt\]\[\int\limits_{ }^{ }(t+2)^2\sqrt{t+2}-(4t+4)\sqrt{t+2}dt\]\[\int\limits_{ }^{ }(t+2)^{5/2}-(4t+4)\sqrt{t+2}dt\]

OpenStudy (idku):

Doing the magic zero again. \[\int\limits_{ }^{ }(t+2)^{5/2}-(4t+4)\sqrt{t+2}-2\sqrt{t+2}+2\sqrt{t+2}~dt\]\[\int\limits_{ }^{ }(t+2)^{5/2}-(4t+4+2)\sqrt{t+2}+2\sqrt{t+2}~dt\]\[\int\limits_{ }^{ }(t+2)^{5/2}-(4t+8)\sqrt{t+2}+2\sqrt{t+2}~dt\]\[\int\limits_{ }^{ }(t+2)^{5/2}-4(t+2)\sqrt{t+2}+2\sqrt{t+2}~dt\]

OpenStudy (idku):

\[\int\limits_{ }^{ }(t+2)^{5/2}-4(t+2)^{3/2}+2\sqrt{t+2}~dt\]

OpenStudy (p0sitr0n):

interesting trick tho

OpenStudy (idku):

So if you haven't learned integration by parts, this is pretty much the same thing, but you would be doing Algebra instead of integration. you are still increasing the exponents of (t+2) as you would, when doing integration by parts.

OpenStudy (idku):

@gbblue you should be able to take it from here.

OpenStudy (idku):

lol

OpenStudy (idku):

it is just like integrating csc(x), but substituting 2u instead of x, and then rewriting the bottom with sin(a+a) rule, and on....

OpenStudy (idku):

@gbblue do you feel comfortable with any of the posted approaches? Or do you want to do integration by parts instead. or do you just want to get a bad grade?

OpenStudy (anonymous):

Lol.@idku for you last comment.:P

OpenStudy (idku):

I wanted to leave the last comment, or second to last to gbblue , but he/she doesn't seem to want to do the problem.

OpenStudy (idku):

I am not coming back here again....

OpenStudy (anonymous):

sorry I was getting coffee, but yes this was very helpful and I can take it from here

OpenStudy (anonymous):

@gbblue check my method too

OpenStudy (idku):

Finally.... it's alright. I hope you have a good coffee for the least oart.

OpenStudy (anonymous):

@Princer_Jones I do not quite understand your method

OpenStudy (anonymous):

it is integration by trigonometric substitution